5. Calculate the lipids required in grams for a 37 year old male patient weighing 240 pounds and 6 feet tall. This patient has suffered from severe trauma. (Note: Use the lower limit of fat range & a 10% lipid emulsion)

PLEASE MAY I HAVE AN ANSWER STEPBY STEP
Answer #1

The lipids required in grams for a 37 year old male patient weighing 240 pounds and 6 feet tall. This patient has suffered from severe trauma. (Note: Use the lower limit of fat range & a 10% lipid emulsion)

For this we have t calculate first the daily energy requirement

BMR= 655 + ( 4.35 x weight in pounds ) + ( 4.7 x height in inches ) - ( 4.7 x age in years )

( 4.35 x weight in pounds )

= 4.35 x 240

= 1044

( 4.7 x height in inches )

= 4.7 x 72

= 338.4

= ( 4.7 x age in years )

= 4.7 x 37

= 173.9

= 655 + 1044 + 338.4 -173.9

= 2037.4 -173.9

= 1863.5

Calories required per day = BMR x Activity factor

= 1863.5 x 1.2

= 2236.2 Calories

Out of this; 30% energy should be from fat

= 2236.2 x .30

= 670.86 should come from fat

= 0.670 KCalories

= Fat emulsions: ○ 10% fat emulsion: 1.1 kcal/cc

= 1.1 kcal/ gram

= 0.670/ 1.1

= 0.609 grams